OpenStudy (anonymous):

i need to solve this problem using the multiplication principle (2/7) = (Z/-3) and check. I got the answer z=-6/7 how do I check it?

OpenStudy (anonymous):

substitute it back it. 2/7 = (-6/7)/-3 -6/7 / -3 = -6/7 * -1/3 =6/21 = 2/7